Villanueva Corp. had 200,000 shares of common stock, 20,000 shares of convertible preferred stock, and $1,000,000 of 5% convertible bonds outstanding during 2020. The preferred stock is convertible into 40,000 shares of common stock. During 2020, Villanueva paid dividends of $1.20 per share on the common stock and $4 per share on the preferred stock. The bonds were originally issued at par and each $1,000 bond is convertible into 45 shares of common stock. Villanueva’s net income for 2020 was $1,800,000 and the applicable income tax rate was 25%. Compute Villanueva's Basic EPS for 2020.

A. $8.6 per share

B. $9.0 per share

C. $9.4 per share

D. None of above

Refer to the information above. Compute Villanueva's Diluted EPS for 2020.

A. $6.04 per share

B. $6.45 per share

C. $6.49 per share

D. $6.60 per share
Answer #1

Basic EPS = (Net income - Preferred stock dividends) / Weighted average number of common stock outstanding

= [$1,800,000 - (20,000*$4)] / 200,000

= $8.6 per share

---------------------------------------------------------------

Bond interest = $1,000,000 * 5% = $50,000

Tax saving = $50,000*25% = $12,500

Increase in Net income = $50,000 - $12,500 = $37,500

Bonds convertible to common stock = $1,000,000 / $1,000 * 45 = 45,000

Preferred stock convertible to common stock = 40,000

Dilutes EPS = Net income / Shares outstanding

= ($1,800,000+$37,500) / (200,000+45,000+40,000)

= $6.45 per share